a dairy farmer plans to enclose a rectangular pasture adjacent to a river. the pasture must contain 180,000 square meters. no fencing is required along the river. what dimensions will use the smallest amount of fencing?
so confusing?? please help

Answer by asha(30)

let length be x metres and breadth be y metres
given Area of rectangular pasture =180,000square metres
x*y =180,000
also x+2y should be the least for the values of x and y
factors of 18 are 6*3 ,18*1 and 9*2
if x=6and y=3 then x+2y=12
ifx=18 and y=1 then x+2y=20
if x=9 and y=2 then x+2y=13
x=6 and y=3 give us the least value of x+2y
so the cost of fencing is the least if lenght is 600ms. and breadth is 300 ms.
good luck !!!